40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Mar 28 03:37:07 2013, Zach, Optics, Configuration, Test setup input optics progress OMC_test_setup_IO.jpg2013-03-27_05.48.20.jpg2013-03-28_02.29.36.jpg
    Reply  Fri Mar 29 08:55:00 2013, Zach, Optics, Configuration, Beam launched into fiber 
       Reply  Thu Apr 4 23:44:52 2013, Koji, Optics, Configuration, Beam launched into fiber 
          Reply  Fri Apr 5 18:18:36 2013, Zach, Optics, Configuration, AOM probably broken Isomet_1250C_zin_4_5_13.png
Message ID: 97     Entry time: Thu Apr 4 23:44:52 2013     In reply to: 87     Reply to this: 99
Author: Koji 
Type: Optics 
Category: Configuration 
Subject: Beam launched into fiber 

We had to move our flipper mirror to share the beam between Peter's setup and ours as our flipper is at the place where the ISS PD array base is supposed to be!
There was no place to insert the flipper in the setup. We (Peter and Koji) decided to move the laser back for ~2".

This entirely changed the alignment of the setup. The fiber coupler was my reference of the alignment.
Once the beam is aligned, I check the coupling to the fiber. It was 50%.

I tweaked the lens and eventually the coupling is improved to 83%. (24.7mW incident, 20.4mW obtained.)

Then, I started to check the AOM path. I noticed that the 1st (or -1st) order beam is very weak.
The deflection efficiency is ~0.1%. Something is wrong.
I checked the driver. The driver's coupler output (1:10) show the amplitude ~1V. (good)
I check the main output by reducing the offset. When the coupler output is 100mV, the main output was 1V. (good)
So is the AOM itself broken???

ELOG V3.1.3-